“Killing Season” is a British medical term used to describe the period in August when newly qualified doctors enter the National Health Service. Around this time there is an up tick in medical misdiagnosis and surgical complications. A similar phenomenon exists in the US and occurs in the month of July3 . The probability that a doctor correctly diagnoses a particular illness is 0.7. Given that the doctor makes an incorrect diagnosis, the probability that the patient files a lawsuit is 0.4. At any given time 15% of all doctors are facing lawsuits.

(a) What is the probability that a doctor facing a lawsuit made a correct diagnosis.

(b) What is the probability that a doctor facing a lawsuit made an incorrect diagnosis?
Answer #1

a)P(correct diagnosis|lawsuit )=P(correct diagnosis and lawsuit)/P(lawsuit) =(0.15-(1-0.7)*0.4)/0.15 =0.2

b) P(Incorrect diagnosis|lawsuit )=P(incorrect diagnosis and lawsuit)/P(lawsuit) =(1-0.7)*0.4/0.15 =0.8